OpenStudy (anonymous):

I really need help with trig. Just a few problems >.< *Will give medal and fan*

OpenStudy (anonymous):

tan (sin^-1 (-1))=?

OpenStudy (tkhunny):

The sine of what angle produces -1?

OpenStudy (anonymous):

3pi/2, right?

OpenStudy (anonymous):

I don't understand, The multiple choice answers say 1, 0. -1 or undefined as the answers.

OpenStudy (dumbcow):

yes its 3pi/2 now what is tan(3pi/2) ?

OpenStudy (anonymous):

0/-1? which is undefined, right?

OpenStudy (dumbcow):

yep :)

OpenStudy (anonymous):

oh! okay, thanks!

OpenStudy (tkhunny):

Well, normally, the inverse sine is defined on \([-\pi/2,\pi/2]\), so we should have been talking about \(-\pi/2\) rather than \(3\pi/2\). The result is the same.